A boat is traveling at 4.8 m/s in the same direction as an ocean wave of wavelength 30 m and speed 6.8 m/s. If the boat is on the crest of a wave, how much time will elapse until the boat is next on a crest?
Answer #1

The relative speed of the boat is, V, = V2 - v = 6.8 m/s - 4.8 m/s=2 m/s The rate of change in distance is called as velocity